Say mr karim runs on a plain land and jumps from a spot in forward direction. Recognize and describe his jumping motion through an equation and make an argument for your answer.


1
Expert's answer
2021-10-31T18:08:43-0400

Motion of Mr Karim is projectile motion, whose equation is obtained as follows. Let his initial velocity just before jump be v0 and angle by θ. Thus his initial velocity has two components:

horizontal or x-component "=v_0 cos \u03b8"

vertical or y-component "= v_0 sin \u03b8"

Let (x,y) be his position at any time t, thus

"x = v_0 cos θ t \\

t = \frac{x}{v_0 cos θ} \\

y = v_0 sin θ t - \frac{1}{2}gt^2 \\

= \frac{v_0 sin θx}{v_0 cos θ} - \frac{1}{2}\frac{gx^2}{v^2_0 cos^2 θ} \\

y = (tan θ)x - (\frac{1}{2} \times \frac{g}{v^2_0 cos^2 θ})x^2"

This is the equation representing the trajectory of Mr Karim during his jump.